OpenStudy (anonymous):

A State Park charges $5 per car plus $1 per person as an admission fee. The total charged for a car with x people is f(x) = x + 5. How will the graph of this function change if the per car charge is changed to $8 per car? The line will shift vertically down by $3. The line will shift vertically up by $3. The line will shift vertically up by $8. The line will shift vertically down by $8. @Directrix @Mesopretty

Directrix (directrix):

Hold on a second and let me think.

OpenStudy (anonymous):

ok

Directrix (directrix):

The lines y = x + 5 and y = x + 8 are parallel because they have the same slope. To see what's going on, look at the y-intercepts. With the increase, the y-intercept jumped upward (vertical) $3 from what it was.

Directrix (directrix):

Question?